A 910 gram bullet is fired into a 30.0 kg block of 293 K ballistic gel. The bullet and the block then combine and slide 1.06 m across the floor (uk=0.20) to a stop where the block of gel reaches a temperature of 302 K, Estimate the speed of the projectile before it hit the block? Please show work!!
Answer #1

KE = work done by friction + heat

0.5*(m1+m2)*V^2 = u*N*d + msdT

specific heat of ballistic gel = 1.439 x10^3 J/Kg K

0.5*(30+0.91)*V^2 = 0.2*(30+0.91)*9.8*1.06 + (30+0.91)*1.439 x10^3*(302-293)

V^2 = 160.97 m/s

conservation of momentum

m1u1 + m2u2 = (m1+m2)*V

0.91*u1 + 0 = (30+0.91)*160.97

u1 = 5467.68 m/s